# Near-Field THz Bending Beamforming: A Convex Optimization Perspective

### v1

@@ I Introduction
Fortunately, recent advances in wavefront engineering have demonstrated the potential of non-diffracting beams for achieving reliable wireless communications even with obstacles in the LoS direction, by exploiting their inherent “self-healing” properties. For example, one typical non-diffracting beam is referred to as bending beam, which is able to focus the beam energy on a prescribed curved trajectory in the non-LoS direction, as illustrated in Fig. 1. By this means, a base stations (BS)/ access point (AP) can achieve consistently reliable communications with dynamically blocked users, thus significantly enhancing the THz signal coverage, especially in the complex environment with dense obstacles. The fundamental physics of such bending beams can be characterized through the stationary solutions to the celebrated Helmholtz equation, which governs the spatial evolution of EM wavefronts[8, 9, 10]. A pioneering work [11] has demonstrated THz bending beam using an ELAA system. However, this method can only generate a specific type of trajectory for the bending beam, i.e. Airy beam, by following the Airy function, which may limit its applications in practice. To enable more flexible bending beamforming, a recent work [12] proposed an alternative method by assuming a continuous antenna aperture and determining the phase profile at any point based on its tangent line to the desired trajectory. However, this method may still achieve suboptimal performance in general, due to its reliance on the geometric properties of the desired trajectory and other limitations, as will be detailed in this paper later.
@@
To address the above challenges, we propose a new and general bending beamforming method by adopting the convex optimization techniques. Unlike the physics-driven approaches as in [11] and [12], we formulate the bending beamforming design as an optimization problem. Specifically, we aim to maximize the minimun received signal power among all positions along the desired propagation trajectory by optimizing the digital or analog beamformer at the BS. However, the resulting problem is non-convex, making it difficult to be solved optimally. To address this difficulty, we employ the successive convex approximation (SCA) technique to obtain a high-quality suboptimal solution by solving a series of approximate convex optimization problems iteratively. Numerical results demonstrate that our proposed bending beamforming method outperforms the traditional method and shows robustness to the obstacles in the environment.
@@
Notations: $\mathbb{R}$ and $\mathbb{C}$ represent the sets of real and complex numbers, respectively. For a complex number $a$, $\operatorname{Re}\{a\}$ and $|a|$ denote its real part and modulus, respectively. For a complex-valued vector $\mathbf{x}$, $\mathbf{x}^{\text{H}}$, $\|\mathbf{x}\|_{2}$, and $x(n)$ denote its transpose, conjugate transpose, $\ell_{2}$-norm, and the $n$-th element, respectively. For a matrix $\bm{A}$, $\operatorname{Tr}(\bm{A})$ represents its trace. $f^{\prime}(x)$ denotes the first-oder derivative of $f(x)$.
@@ V Numerical Results
In this section, we present numerical results to evaluate the performance of the proposed method for generating bending beams. Unless otherwise stated, the simulation parameters are set as follows. The carrier frequency is $f_{c}=300$ GHz. The number of antennas at BS is $N=400$ and the spacing between any two adjacent antennas is set to half-wavelength. As such, the Rayleigh distance can be calculated as 80 m. The user’s position is located at $(x,z)=(225\beta,15)$. The antenna array is assumed to be located in the $x$-axis, and the beam trajectory is parabolic-shaped and given by $x=\beta z^{2},0\leq z\leq 15$. The number of sampling points along the bending beam trajectory is set to $M=200$. For performance comparision, we show the performance achieved by the proposed scheme with ABF and DBF, as well as the traditional tangent method (TM) in [12].
@@
Fig. 3 plots the bending beams generated by different schemes for different values of $\beta$, where the desired beam trajectory is indicated by the blue dashed lines. It is observed that for each $\beta$ considered, the proposed scheme can generate high received signal powers in the vicinity of the beam trajectory with both ABF and DBF. This observation confirms that the proposed bending beamforming method exhibits the characteristics similar to the non-diffracting beam.
@@
To gain more insights, we plot in Fig. 4 the received signal power at all positions along the desired trajectory. It is observed that the proposed max-min bending beamforming method yields an approximately uniform distribution of the received signal power along the desired trajectory with either ABF or DBF, resulting in a higher max-min received signal power and the received signal power at the user compared to the TM. It is also observed that the DBF achieves a similar performance to that of the ABF, which contradicts our initial expectations. Since the DBF offers more degrees of freedom in design, we anticipated superior performance. However, this result aligns with prior wavefront engineering studies cited earlier in this paper, where phase control is shown to play a significant role in generating curved beams.
In the previous examples, we show the performance of the proposed method without actual obstacles in the environment. To evaluate the robustness of the proposed method in the presence of obstacles, we show in Fig. 5 its generated bending beam under the ABF, where an obstacle is placed between the antenna array and the user. If the LoS channel between the $n$-th antenna and any position $(x,z)$ is blocked by the obstacle, we set $h_{n}(x,z)=0$. The desired beam trajectory is set to $x=0.0008z^{2},0\leq z\leq 15$m. As observed from Fig. 5, even the obstacle can nullify the LoS channels between several antennas and positions, the generated bending beam is able to maintain a pre-defined shape without significant deviation, showing a high robustness to the presence of the obstacle. It thus follows that the proposed max-min bending beamforming method can provide an efficient solution to resolve the challenging issue of THz signal blockage.
Figure 5: Bending beam generated by the proposed method with a blockage.
@@ VI Conclusion
In this paper, we proposed a new and general bending beamforming method by formulating a max-min beamforming approach. Unlike the conventional tangent method, our proposed method directly tackles the actual received signal power and significantly reduces the reliance the geometric properties of the desired trajectory. To solve the non-convex max-min optimization problem in the proposed method, we adopt the SCA algorithm to obtain a locally optimal solution. Simulation results demonstrate that the proposed method achieves superior performance to the tangent method in terms of the max-min received signal power and the received signal power at the user.

### v2

@@ I Introduction
Fortunately, recent advances in wavefront engineering have demonstrated the potential of non-diffracting beams for achieving reliable wireless communications even with obstacles in the LoS direction, by exploiting their inherent “self-healing” properties. For example, one typical non-diffracting beam is referred to as bending beam, which is able to focus the beam energy on a prescribed curved trajectory in the non-LoS direction, as illustrated in Fig. 1. By this means, a base stations (BS)/ access point (AP) can achieve consistently reliable communications with dynamically blocked users, thus significantly enhancing the THz signal coverage, especially in the complex environment with dense obstacles. The fundamental physics of such bending beams can be characterized through the stationary solutions to the celebrated Helmholtz equation, which governs the spatial evolution of EM wavefront[8, 9, 10]. A pioneering work [11] has demonstrated THz bending beam using an ELAA system. However, this method can only generate a specific type of trajectory for the bending beam, i.e. Airy beam, by following the Airy function, which may limit its applications in practice. To enable more flexible bending beamforming, a recent work [12] proposed an alternative method by assuming a continuous antenna aperture and determining the phase profile at any point based on its tangent line to the desired trajectory. However, this method may still achieve suboptimal performance in general, due to its reliance on the geometric properties of the desired trajectory and other limitations, as will be detailed in this paper later.
@@
To address the above challenges, we propose a new and general bending beamforming method by adopting the convex optimization techniques. Unlike the physics-driven approaches as in [11] and [12], we formulate the bending beamforming design as an optimization problem. Specifically, we aim to maximize the minimum received signal power among all positions along the desired propagation trajectory by optimizing the digital or analog beamformer at the BS. However, the resulting problem is non-convex, making it difficult to be solved optimally. To address this difficulty, we employ the successive convex approximation (SCA) technique to obtain a high-quality suboptimal solution by solving a series of approximate convex optimization problems iteratively. Numerical results demonstrate that our proposed bending beamforming method outperforms the traditional method and shows robustness to the obstacles in the environment.
@@
Notations: $\mathbb{R}$ and $\mathbb{C}$ represent the sets of real and complex numbers, respectively. For a complex number $a$, $\operatorname{Re}\{a\}$ and $|a|$ denote its real part and modulus, respectively. For a complex-valued vector $\mathbf{x}$, $\mathbf{x}^{\text{H}}$, $\|\mathbf{x}\|_{2}$, and $x(n)$ denote its transpose, conjugate transpose, $\ell_{2}$-norm, and the $n$-th element, respectively. For a matrix $\bm{A}$, $\operatorname{Tr}(\bm{A})$ represents its trace. $f^{\prime}(x)$ denotes the first-order derivative of $f(x)$.
@@ V Numerical Results
In this section, we present numerical results to evaluate the performance of the proposed method for generating bending beams. Unless otherwise stated, the simulation parameters are set as follows. The carrier frequency is $f_{c}=300$ GHz. The number of antennas at the BS is $N=400$ and the spacing between any two adjacent antennas is set to half-wavelength. As such, the Rayleigh distance can be calculated as 80 m. The user’s position is located at $(x,z)=(225\beta,15)$. The antenna array is assumed to be located in the $x$-axis, and the beam trajectory is parabolic-shaped and given by $x=\beta z^{2},0\leq z\leq 15$. The number of sampling points along the bending beam trajectory is set to $M=200$. For performance comparision, we show the performance achieved by the proposed scheme under ABF and DBF, as well as the traditional tangent method (TM) in [12].
@@
First, Fig. 3 plots the bending beams generated by different schemes for different values of $\beta$, where the desired beam trajectory is indicated by blue dashed lines. It is observed that for each $\beta$ considered, the proposed scheme yields high received signal power along the desired trajectory for both ABF and DBF. This result suggests that the beam generated by the proposed method exhibits non-diffracting characteristics similar to that by the conventional TM, as shown in Fig. 3(c), Fig. 3(f), Fig. 3(i).
@@
To provide a quantitative performance comparison between the proposed method and the TM, we plot in Fig. 4 the received signal power at all positions along the desired trajectory by different schemes. It is observed that compared to the TM, the proposed max-min bending beamforming method achieves a more uniform distribution of the received signal power along the desired trajectory for both ABF and DBF, resulting in a higher max-min received signal power and received signal power at the user. It is also observed that the ABF achieves a comparable performance to the DBF. This observation aligns with the prior studies on wavefront engineering in [12], which highlight the dominant role of phase control over amplitude control in generating bending beams. It is also observed that the performance gap between the proposed method and the TM increases with $\beta$, suggesting that the proposed method is more preferred for trajectories with large curvature compared to the TM. This is consistent with our discussion at the end of Section III.
In the previous examples, we show the performance of the proposed method without accounting for actual obstacles in the environment. To evaluate its robustness in the presence of obstacles, we show in Fig. 5 the bending beam generated under ABF, where an obstacle is placed between the antenna array and the user. If the LoS channel between the $n$-th antenna and any position $(x,z)$ is blocked by the obstacle, we set $h_{n}(x,z)=0$. The desired beam trajectory is set to $x=0.0008z^{2},0\leq z\leq 15$m. As observed from Fig. 5, even the obstacle can nullify the LoS channels between several antennas and positions, the generated bending beam is able to maintain a pre-defined shape without significant deviation, showing high robustness against obstructions. It thus follows that the proposed max-min bending beamforming method offers an efficient solution to resolve the challenging issue of THz signal blockage.
Figure 5: Bending beam generated by the proposed method in the presence of a blockage.
@@ VI Conclusion
In this paper, we proposed a new and general bending beamforming method by formulating a max-min beamforming approach. Unlike the conventional tangent method, our proposed method directly tackles the actual received signal power and significantly reduces the reliance the geometric properties of the desired trajectory. To solve the non-convex max-min optimization problem in the proposed method, we adopted the SCA algorithm to obtain a locally optimal solution. Simulation results demonstrated that the proposed method achieves superior performance over the tangent method in terms of the max-min received signal power and the received signal power at the user.